OpenStudy (anonymous):

Multiply, using any appropriate method. (3-a)^3

OpenStudy (anonymous):

Writing this out and then FOILing might helping you. this can be rewritten as \[(3-a) * (3-a) * (3-a)\] Then just FOIL the first two together so (3-a) * (3-a), after you've found that, multiply it by the last (3-a). Are you familiar with FOIL? (First, Outer, Inner, Last)

OpenStudy (anonymous):

I am familiar with it but it gets me really lost

OpenStudy (anonymous):

ok, to FOIL (3-a) * (3-a) F(irst) - multiply the first term in each expression: 3*3... so we have 9. O(uter) - we multiply the terms that are on the outside edges. That would be 3 * -a, which is just -3a. I(nner) - we multiply the two inside terms. -a*3. Another -3a. L(ast) - multiply the last term in each expression. a*a = a^2

OpenStudy (anonymous):

Now put those terms together. we have a 9 - 3a - 3a - a^2. Combine the 3a's and you've got \[9 - 6a - a^2\]

OpenStudy (anonymous):

Now just multiply that by the remaining (3-a) , you will have to combine a few terms, then your done!

OpenStudy (whpalmer4):

Do it one step at a time. You've got \[(3-a)^3=(3-a)(3-a)(3-a)\]Tackle just a pair of \((3-a)\)'s first, then multiply the result by another \((3-a)\) for the whole thing. \[(3-a)(3-a) = 3(3-a) - a(3-a)\]by distributive property Now do that again \[3(3-a)-a(3-a) = 3*3-3*a-a*3-a*(-a) \]\[= 9-3a-3a+a^2\]\[ = 9-6a+a^2\] Now multiply that by \((3-a)\) to get \((3-a)^3\) \[(3-a)(9-6a+a^2) = 3(9-6a+a^2) -a(9-6a+a^2) \]\[= 3*9-3*6a+3*a^2-a*9-a(-6a)-a*a^2 \]\[= 27-18a + 3a^2-9a+6a^2-a^3\]\[=-a^3+9a^2-27a+27\]

OpenStudy (anonymous):

ok now I am really lost because I did what you said and when I put it in it said that it was wrong the correct answer was \[1-3a+3a ^{2}-a ^{3}\]

OpenStudy (anonymous):

how did they come up with that answer. I am so lost on trying to figure this out

OpenStudy (anonymous):

Im not sure, while I don't think its a good idea to give an answer, the answer whpalmer4 gave you is 100% correct.

OpenStudy (anonymous):

ok can you walk me through doing \[(4-3a)^{3}\]

OpenStudy (anonymous):

Was there something missing from what you gave us to begin with? Or was the problem simply \[3-a)^3\]

OpenStudy (anonymous):

this is the same basic idea as above. If you replace the 3 with a 4, and the "a" with a "3a" and follow the steps, you will get the right answer.

OpenStudy (whpalmer4):

The problem for which \[1-3a+3a^2-a^3\] is the correct answer is \((1-a)^3\)

OpenStudy (anonymous):

write it as \[(4-3a) * (4-3a) * (4-3a)\] and just work on a piece at a time.

OpenStudy (anonymous):

ok this is what I have for the first two steps

OpenStudy (whpalmer4):

@xartaan as for giving the answer, I know she's got a ton of these to do, so she's not going to get any appreciable advantage from my having worked one through to the end, except seeing how to do it.

OpenStudy (anonymous):

\[(4-3a)^{3}=(4-3a)(4-3a)(4-3a)\] \[(4-3a)(4-3a)=4(4-3a)-a(4-3a)\]

OpenStudy (whpalmer4):

@xartaan the code of conduct says we are to guide them — she's being guided through the process. if she was just an answer seeker, she wouldn't have hung in there this long!

OpenStudy (whpalmer4):

@cahayes9498 that's very close, except you turned the 3a into a at the end... \[(4-3a)(4-3a) = 4(4-3a)-3a(4-3a)\] see the difference?

OpenStudy (anonymous):

then \[4(4-3a)-3a(4-3a)=4*4-4*3a-3a*4-3a*(-3a)\]

OpenStudy (anonymous):

yeah I noticed that after wards I corrected it

OpenStudy (whpalmer4):

good eyes!

OpenStudy (whpalmer4):

looks good so far, now simplify

OpenStudy (anonymous):

now \[4*4-4*3a-3a*4-3a*(-3a)=16-12a-12+3a ^{2}\]

OpenStudy (whpalmer4):

whoops, another typo in there...

OpenStudy (whpalmer4):

only one constant term per multiplication!

OpenStudy (anonymous):

huh? :(

OpenStudy (anonymous):

\[16-12a-12+3a ^{2}=16-24a+3a ^{2}\]

OpenStudy (anonymous):

\[=4(16-24a+3a ^{2})-3a(16-24a+3a ^{2})\]

OpenStudy (anonymous):

ok this is where I start to really get confused

OpenStudy (anonymous):

\[4*16+4*24a+4*3a ^{2}-3a ^{2}*16-3a ^{2}(-24a)-3a*3a ^{2}\] is the next thing that I got

OpenStudy (anonymous):

@whpalmer4 you still there

OpenStudy (whpalmer4):

I am now! :-)

OpenStudy (anonymous):

ok great

OpenStudy (anonymous):

am I correct so far?

OpenStudy (whpalmer4):

\[16-12a-12+3a^2 \ne 16-24a+3a^2\]but \[16-12a-12a+3a^3=16-24a+3a^2\]

OpenStudy (whpalmer4):

and the latter is what you wanted, you just dropped an a again

OpenStudy (whpalmer4):

have to be fastidious when doing this work!

OpenStudy (whpalmer4):

so now you have \((4-3a)^2 = 16-24a+3a^2\) and you have to multiply that by one more \((4-3a)\) in order to get \((4-3a)^3\) \[(4-3a)(16-24a+3a^2)=\] your turn :-)

OpenStudy (whpalmer4):

oh, and I see you already did that...now can you simplify?

OpenStudy (whpalmer4):

except you lost a minus sign in front of -24a when you multiplied :-(

OpenStudy (anonymous):

so it would be \[4*16+4*24a+4*3a ^{2}-3a ^{2}*16-3a ^{2}*24a-3a+3a ^{2}\]

OpenStudy (anonymous):

which would simplify to \[64+96a+12a ^{2}-48a ^{2}-72a ^{2}-3a+3a ^{2}\]

OpenStudy (anonymous):

this is the part that I get so lost with and dont understand how to simplify it the correct way

OpenStudy (whpalmer4):

okay, the problem is you aren't doing the multiplication correctly, so it doesn't matter if you can simplify correctly :-)

OpenStudy (whpalmer4):

Let's look at just the first half: \[(4-3a)(16-24a+3a^2) = 4(16-24a+3a^2) + (\text{the rest})\]\[4(16-24a+3a^2) = 4*16-4*24a + 4*3a^2 + (\text{the rest})\] Look carefully at what I have there vs. what you have there...

OpenStudy (anonymous):

ok I see that I added when I should have subtracted

OpenStudy (whpalmer4):

Yeah, you didn't apply the - negative sign when multiplying 4 * -24a...

OpenStudy (anonymous):

ok so from here I would do 4*16 which is 64

OpenStudy (whpalmer4):

you made the corresponding mistake in the other half, too, it appears...

OpenStudy (whpalmer4):

why don't you write out the last line for me again, the practice would be good...

OpenStudy (anonymous):

\[4*16+4*24a+4*3a ^{2}-3a ^{2}*16+3a ^{2}*24a-3a+3a ^{2}\]

OpenStudy (anonymous):

is that correct?

OpenStudy (whpalmer4):

no, you keep forgetting to include the negative signs! \[4(16-24a+3a^2) = 4*16 + 4*(-24a) + 4*3a^2 = 64-96a+12a^2\]

OpenStudy (anonymous):

in the second part?

OpenStudy (whpalmer4):

no, in the first part. you always write \[4*16+4*24a+4*3a^2\]but it is \[4*16\color{red}{-}4*24a+4+3a^2\]

OpenStudy (anonymous):

I am never going to get this

OpenStudy (anonymous):

well thank you for trying to help me I am just going to get this one wrong I am never going to figure this out I get so confused with it.

OpenStudy (whpalmer4):

no, there's no reason to get this wrong. \[4(x-a) = 4*x + 4*(-a) = 4x-4a\] \[4(x+a) = 4*x + 4*a = 4x+4a\]

OpenStudy (whpalmer4):

you're getting the trickier parts right, and getting tripped up an a relatively easy part :-)

OpenStudy (whpalmer4):

\[4(16-24a+3a^2) = 4*16 + 4*(-24a) + 4*3a^2\]\[=64-96a+12a^2\]

OpenStudy (whpalmer4):

you just can't drop that - sign in front of 24a...

OpenStudy (anonymous):

I just dont understand why it cant be \[(4-3a)^{3} = (1a)^{3}\]

OpenStudy (whpalmer4):

the sign is part of the number. maybe it would help to write it out whatever sign it is: \[4(16-24a+3a^2) = 4*(+16)+ 4*(-24a) + 4(+3a^2) = 64 + (-96a) + 12a^2 \]\[=64-96a+12a^2\]

OpenStudy (anonymous):

ok I understand that part now but the rest of it all is making my head explode

OpenStudy (whpalmer4):

because \(4-3a \ne 1a\) for most values of \(a\). What if \(a=0\)? \[4-3(0) = 1*0\]\[4=0\]Uh, no.

OpenStudy (anonymous):

I know that it dont equal that I was just wishing it would be that simple

OpenStudy (whpalmer4):

Think about numbers. 169 = 1*100 + 6*10 + 9*1, right? well, this is the same thing, except instead of powers of 10, we have powers of \(a\).

OpenStudy (whpalmer4):

\[4-3a\] is really \[-3a + 4 = -3a^1+4a^0\]because \(a^1=a\) and \(a^0=1\) (so long as \(a\ne 0\), as \(0^0\)is undefined)

OpenStudy (anonymous):

ok i got the first part of it with the \[64-96a+12a ^{2}\] now the rest of it that I came up with which is probably wrong is \[-16a+24a ^{2}-3a ^{3}\] so all together I got \[64-96a+12a ^{2}-16a+24a ^{2}-3a ^{3}\]

OpenStudy (whpalmer4):

let's see\[-3a(16-24a+3a^2) = -48a+72a^2-9a^3\]looks like you left off the 3...

OpenStudy (whpalmer4):

all together you should have \[64-96a+12a^2 -48a +72a^2-9a^3\]Can you simplify that by collecting like terms?

OpenStudy (anonymous):

ok so the second part I just had to multiply everything by the 3a?

OpenStudy (whpalmer4):

by the -3a...

OpenStudy (anonymous):

oh ok that makes sense now

OpenStudy (anonymous):

so now I would just combine like terms

OpenStudy (anonymous):

would it be \[64-144a+84a ^{2}-9a ^{3}\]

OpenStudy (whpalmer4):

crap, we've been carrying an error all along, I think. let's go from the start: \[(4-3a)(4-3a) = 4(4-3a) -3a(4-3a) = 4*4-4*3a-3a*4+9a^2 \]\[= 16-24a+9a^2\] Now we multiply again by \((4-3a)\): \[(4-3a)(16-24a+9a^2) = 4(16-24a+9a^2)-3a(16-24a+9a^2)\]\[= 4*16-4*24a+4*9a^2 -3a*16 -3a*(-24a)-3a*9a^2\]\[=64-96a+36a^2-48a+72a^2-27a^3\]

OpenStudy (whpalmer4):

you correctly simplified an incorrect equation :-) good job by you, poor job by me

OpenStudy (whpalmer4):

This last one is correct, I promise!

OpenStudy (anonymous):

so what I got was not right

OpenStudy (whpalmer4):

you simplified what you had in front of you correctly. I'm not sure whose fault it is that what you had to simplify was not correct. I should have spotted the mistake earlier

OpenStudy (anonymous):

\[64-144a+108a ^{2}-27a ^{3}\]

OpenStudy (whpalmer4):

yes!!!!!! :-)

OpenStudy (anonymous):

how about that, with the correction made

OpenStudy (anonymous):

go me.... its my birthday.... LOL

OpenStudy (whpalmer4):

you deserve a medal for hanging in there :-)

OpenStudy (anonymous):

and i have the subtract and add correct?

OpenStudy (whpalmer4):

happy birthday to you, happy birthday to you...how old? double digits yet? :-)

OpenStudy (anonymous):

not really my birthday just a saying in florida sorry

OpenStudy (whpalmer4):

that answer is exactly correct.

OpenStudy (anonymous):

ok good, i am glad I didnt quit

OpenStudy (whpalmer4):

hey, no one is checking IDs here, if you say it's your birthday, it's your birthday :-)

OpenStudy (anonymous):

thank you for being so patient with me and helping me to figure this out

OpenStudy (whpalmer4):

free drinks on the house! help yourself to anything in the refrigerator :-)

OpenStudy (anonymous):

lol.... I am going to help myself to the bed.... Im too old should have been in bed a long time ago

OpenStudy (whpalmer4):

you're welcome! this is the sort of thing where it's puzzling until you have that little breakthrough, and then you wonder why it was such a big problem...